OpenStudy (austinlr12):

Given: AB and CD If the coordinates of point A are (8, 0) and the coordinates of point B are (3, 7), the y-intercept of AB is ________. If the coordinates of point D are (5, 5), the equation of line CD is y = _____ x + _____.

OpenStudy (austinlr12):

@AdoNine @Michele_Laino

OpenStudy (michele_laino):

here we have to write the equation of the line which passes at points A and B, first

OpenStudy (austinlr12):

Okay

OpenStudy (michele_laino):

the slope \(m\) of such line, is given by the subsequent computation: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{7 - 0}}{{3 - 8}} = ...?\]

OpenStudy (michele_laino):

please complete

OpenStudy (austinlr12):

7/5?

OpenStudy (michele_laino):

hint: \(3-8=-5\)

OpenStudy (austinlr12):

7/-5?

OpenStudy (michele_laino):

correct!

OpenStudy (austinlr12):

so thats the first part?

OpenStudy (michele_laino):

the requested equation, is: \[\begin{gathered} y - {y_2} = m\left( {x - {x_2}} \right) \hfill \\ \hfill \\ y - 0 = \frac{{ - 7}}{5}\left( {x - 8} \right) \hfill \\ \end{gathered} \] please simplify

OpenStudy (austinlr12):

y=-7/5x+56/5 is correct ?

OpenStudy (michele_laino):

correct! Next the requested \(y-\)intercept, is the value of \(y\), when \(x=0\), so we have: \[y = \frac{{ - 7}}{5} \cdot 0 + \frac{{56}}{5} = ...?\]

OpenStudy (austinlr12):

y=56/5 ?

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

for second part, I think that coordinates of point C are missing

OpenStudy (austinlr12):

So for the last question I would put y=56/5x+...?

OpenStudy (michele_laino):

we need to know the coordinates of point C

OpenStudy (austinlr12):

Or what would you put after x+ ?

OpenStudy (austinlr12):

On the question it does not have any coordinates for c.

OpenStudy (michele_laino):

how point C is defined?

OpenStudy (austinlr12):

https://snag.gy/Spjkzf.jpg

OpenStudy (michele_laino):

line CD is parallel to line AB, namely, the symbol: \[AB \parallel CD\] means line AB is parallel with respect to line CD

OpenStudy (michele_laino):

now, parallel lines have the same slope, so the slope of line CD is \(m=-7/5\)

OpenStudy (michele_laino):

the requested equation, is therefore: \[y - 5 = \frac{{ - 7}}{5}\left( {x - 5} \right)\] please simplify

OpenStudy (michele_laino):

hint: if we apply the distributive property, we get: \[y - 5 = \frac{{ - 7x}}{5} + 7\]

OpenStudy (austinlr12):

And when I simp. that what I plug into the last part?

OpenStudy (austinlr12):

y=-7/5x+12?

OpenStudy (michele_laino):

yes! Please we have this step: I add \(5\) to both sides, so I get: \[y - 5 + 5 = \frac{{ - 7x}}{5} + 7 + 5\] please simplify

OpenStudy (michele_laino):

correct!